OpenStudy (anonymous):

The sum of two numbers is 48. One number is times as large as the other. What are the numbers?

OpenStudy (unklerhaukus):

your missing something in your question

OpenStudy (anonymous):

The sum of two numbers is what ?

OpenStudy (anonymous):

48/3=16 16 and 2*16=32

OpenStudy (anonymous):

Hmm let one number be n and other 2n Sum: n+2n=3n=48; can you solve it now?

OpenStudy (anonymous):

x + y = 48 (the sum of two numbers is 48) x = 2y (one number is twice as large as the other number) now sub 2y in for x... x + y = 48 2y + y = 48 3y = 48 y = 48/3 y = 16 now sub 16 in for y x = 2y x = 2(16) x = 32 check... x + y = 48 32 + 16 = 48 48 = 48 (correct) your two numbers are 16 and 32